2. A light-bulb manufacturer regularly advertises that his bulbs last 900 hours


with a standard deviation of 75 hours. A random sample is chosen before each


campaign to make sure that the claim is correct. If one such sample of 20 bulbs


show a mean of 925 hours, can the advertising claim be considered an


underestimate at the 0.05 level of significance?


1
Expert's answer
2022-05-13T04:50:11-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le 900"

"H_a:\mu>900"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a right-tailed test is "z_c = 1.6449."

The rejection region for thisright-tailed test is "R = \{z: z >1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{925-900}{75/\sqrt{20}}\approx1.4907"

Since it is observed that "z = 1.4907 <1.6449= z_c ," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=P(Z>1.4907)=0.06802," and since "p=0.06802>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is greater than 900, at the "\alpha = 0.05" significance level.